OpenStudy (anonymous):

Study Guide Help!! 1. What type of information do you need to prove two triangles congruent? 2. Given three coordinates of the vertices of a triangle, how could you prove if it was a right triangle?

OpenStudy (anonymous):

1:IN ORDER to prove two triangles are congruent or not we should know the coordinates of the points on the triangle, 2:through phythagoras theorem that is (hyp)^2=(perp)^2+(base)^2

OpenStudy (anonymous):

So the second one would basically be a^2 + b^2 = c^2

OpenStudy (anonymous):

If the three corresponding vertices of each triangle is given, 1. find the length of each side of the triangles using the distance formula b/n two points in a plane 2. If one side of one triangle is equal in length with one of the sides of the other and it is true for all the three sides, then The two triangles are congruent